A car went 70 miles in 4 hours. If it continues going the same speed how long will it take to go 175 miles?
Found 2 solutions by jorel1380, unlockmath:
Answer by jorel1380(3719)

70/4=175/x
70x=700
x=10
It will take 10 hours.

Answer by unlockmath(1688)

Hello,
Let's set up a proportion as:
70/4=175/x
Cross multiply to get:
70x=700
Divide by 70:
x=10 hours.
Make sense?
RJ
www.math-unlock.com